# EdgeAlpha: Bringing Process Discovery to the Data Sources

### v1

@@ 5.5 Number of Queried Nodes
We now delve deeper into the specifics of the Hospital Log and the Sepsis Cases log. Figure 7 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events. There is still a higher amplitude after 5,000 events. However, there are no long-lasting deviations. This event log serves as a good example of stabilization due to its large activity set. Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
5.5.3 Activity-wise Analysis.
@@
Figure 8 takes a closer look at the specifics of the Sepsis Cases event log. Depending on the activity under examination, the Sepsis Cases log has varied querying patterns. In contrast to other event logs, the Sepsis Cases do not have a designated start event, but six different ones. However, in about 95% of the cases, a case starts with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others. Although ”Release E” shows an increased query count, it is not a start activity. The increased query count is the result of the limited occurrence (only six times), indicating ongoing stabilization. The situation is similar for the four activities to its left. They occur infrequently, which also explains the higher standard error.
5.5.4 Batch Size Analysis.
@@
Figure 6 illustrates the reduction in the average number of queried nodes per event as the batch size increases. The rate in which the number of queries decreases varies between the datasets. However, for all datasets the rate decreases with a growing batch size. For all datasets, the average number of queried nodes per event already lies under 2.5% nodes in the network with a batch size of 40. For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log those are on average 15.6 nodes.
@@
Overall, our results indicate that the reduction in the number of queried nodes depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors of each activity. The experiments show that it takes the activities some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes. While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes. One such adaptation involves employing a sliding window approach, where nodes track activity predecessors within a fixed window. This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
5.6 Footprint Matrix Sizes
@@
In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages. These messages include the node’s local FM along with two additional Booleans indicating if a node recorded the start or end activity of a trace. Thus, the scalability of these messages depends highly on the size of each partial FM. In EdgeAlpha, each node only keeps track of their predecessors in their FMs, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities. Even for datasets like the Hospital Log that has an activity count of 624, each node transmits only a list consisting of at most 624 elements. Compared to existing distributed Alpha Miners [11, 14], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size. Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources. For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
5.7 Evaluation Discussion
@@
EdgeAlpha leverages data locality to enhance its efficiency. Unlike existing distributed Alpha Miners [11, 14] which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source. As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm. Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events. Thus, building a process model becomes substantially faster. This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
6 Conclusion
@@
In this paper, we introduce EdgeAlpha, a novel distributed approach tailored for process discovery in IoT networks. By decentralizing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches. We discuss how the design principles of EdgeAlpha can be applied to other process mining algorithms that construct footprint matrices or record direct successions in the beginning, such as the Heuristics Miner. Moreover, the integration of Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network. The batching of predecessor queries further minimizes communication overhead. Empirical evaluations showcase EdgeAlpha’s immense reduction of messages in real-world scenarios compared to a naive predecessor determination approach. EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event. By batching predecessor queries in groups of 50 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes. We highlight, how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities. By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.
{credits}
6.0.1 Acknowledgements
This work received funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation), grant 496119880.

### v3

@@ V-F2 Stabilization Analysis
We now delve deeper into the specifics of the Hospital Log. Fig. 8 illustrates how the number of queried nodes stabilizes in the Hospital Log after fewer than 3,000 events. There is still a higher amplitude after 5,000 events. However, there are no long-lasting deviations. This event log is a good example of stabilization due to its large activity set. Datasets with more activities take longer to stabilize, as each activity must occur multiple times for the MFPs to be identified.
V-F3 Activity-wise Analysis
@@
Fig. 9 takes a closer look at the specifics of the Sepsis Cases log. Depending on the activity, the Sepsis Cases log shows varied querying patterns. In contrast to other event logs, the Sepsis Cases do not have one designated start event but six different ones. However, about 95% of cases start with ”ER Registration”, which explains why the corresponding bar clearly stands out above the others. Although ”Release E” shows an increased query count, it is not a start activity. The increased query count results from the limited occurrence (only six times), indicating ongoing stabilization. The situation is similar for the four activities to its left. They occur infrequently, which also explains the higher standard error.
V-F4 Batch Size Analysis
@@
Fig. 7 illustrates the reduction in the average number of queried nodes per event as the batch size increases. The rate in which the number of queries decreases varies between the datasets. However, for all datasets the rate decreases with a growing batch size. For all datasets, the average number of queried nodes per event already lies under 2.5% of nodes in the network with a batch size of 40. For the Traffic Fine datasets, 2.5% are on average 0.25 nodes, whereas for the Hospital Log, those are on average 15.6 nodes.
@@
Overall, our results indicate that the reduction in the number of queries depends on the statistical properties of the respective dataset, such as average case lengths, occurrences of self-loops, and the variety of predecessors per activity. The experiments show that it takes the nodes some time to identify their most frequent predecessors and thus to stabilize their average number of queried nodes. While MFP Requesting is suitable for stable processes, adjustments can enhance its suitability for dynamic process changes. One such adaptation involves employing a smaller sliding window, where nodes track activity predecessors only within a fixed window. This strategy leverages recent data for predictions, amplifying the impact of predecessor changes while concurrently minimizing memory usage.
V-G Footprint Matrix Sizes
@@
In the second phase of EdgeAlpha, nodes transmit their data to the central entity, generating a fixed number of messages. These messages include the node’s local FM along with two additional Booleans indicating whether a node recorded the start or end activity of a trace. Thus, the scalability of these messages depends highly on the size of each partial FM. In EdgeAlpha, each node only keeps track of their predecessors in its FM, resulting in transmitting FMs across the network of a maximal size of $n\times 1$, where $n$ represents the total number of activities. Even for datasets like the Hospital Log, which has an activity count of 624, each node transmits only a list consisting of at most 624 elements. Compared to existing distributed Alpha Miners [19, 9], which often require sending up to quadratic FMs to a central entity, our approach is characterized by a compact message size. Moreover, storing such matrix sizes on IoT nodes is feasible even with constrained resources. For long-running processes, EdgeAlpha can further discard event data related to own, predecessor, and successor events once a node’s partial FM has been collected.
V-H Evaluation Discussion
@@
EdgeAlpha leverages data locality to enhance its efficiency. Unlike existing distributed Alpha Miners [19, 9], which typically collect the event data completely before processing it and creating process models, EdgeAlpha adopts a proactive strategy by processing data at runtime and at its source. As soon as a process model is requested, EdgeAlpha starts at a much more advanced point in the Alpha Miner algorithm. Consequently, our strategy provides a significant time advantage, especially when handling large datasets with hundreds of activities and thousands of events. Thus, building a process model becomes substantially faster. This approach is particularly beneficial for iterative real-time process model creation or monitoring tasks.
VI Conclusion
@@
In this paper, we introduce EdgeAlpha, a novel distributed approach tailored to process discovery in IoT networks. By distributing computations to IoT and edge devices, EdgeAlpha addresses scalability and privacy concerns inherent in centralized approaches. We discuss how EdgeAlpha’s design principles can be applied to other process mining algorithms that initially construct footprint matrices or record direct successions, such as the Heuristics Miner [14]. Moreover, integrating Most-Frequent-Predecessor Requesting reduces communication overhead, ensuring efficient event handling within the network. The batching of predecessor queries further minimizes communication overhead. Empirical evaluations showcase EdgeAlpha’s immense reduction of messages in real-world scenarios compared to a naive predecessor determination approach. EdgeAlpha achieves a reduction in communication for determining predecessor events by up to 96% compared to querying all network nodes after each event. By batching predecessor queries in groups of 40 events, experiments show an average reduction in the number of queried nodes per event to less than 2.5% of all nodes. We analytically and experimentally demonstrate the correctness of the algorithm and additionally show how the intermediate results of EdgeAlpha quickly converge to the final result of the Alpha Miner. We highlight how EdgeAlpha provides a significant time advantage over algorithms lacking pre-processing capabilities while still producing the same results as the Alpha Miner. By avoiding the storage of full event traces and sensitive personal data in central locations, EdgeAlpha ensures data privacy through aggregation that eliminates personal identification.